Consider using CSMA/CD for the channel access protocol on a 10 Mbps Ethernet. Assume NODE X has encountered 5 collisions for the same frame:

a) What is the probability that this node chooses K=6 after the 4th collision?

b) If K=6 is the random value chosen, how many seconds would the node delay before sensing the channel again?
Answer #1

a)What is the probability that this node chooses K=6 after the 4th collision?

The probability that a node will choose k=6 after the 4th collision in CSMA/CD is:

25 = 32
p = 1/32
p = 0.03125
p = 0.03125 * 100
p = 3.125%

b) If K=6 is the random value chosen, how many seconds would the node delay before sensing the channel again?

The result of k = 6 corresponds to what delay in seconds on a 10 Mbps Ethernet:
Delay = K * 512 bit times and k = 6
Delay = 6* 512 = 3,072 bit times

3,072bit times/10 mbps=3.072 *10-4 sec.